OpenStudy (anonymous):

x cubed over x squared - 4 divided by x cubed - 9x squared over x squared + 7x -18

OpenStudy (anonymous):

could someone help me get started please

OpenStudy (amistre64):

its difficult to parse the question; can you rewrite it into something more discernable please?

OpenStudy (amistre64):

or use () paranthesis around stuff that is supposed to be grouped together? like perhaps: (x cubed over x squared) - (4 divided by x cubed - 9x) squared over (x squared + 7x) -18

OpenStudy (anonymous):

x cubed over x squared - 4 divided by x cubed - 9x squared over x squared + 7x- 18

OpenStudy (anonymous):

x^3/x^2-4/x^3-9x^2/x^2+7x-18

OpenStudy (anonymous):

Need to know where parentheses are.

OpenStudy (anonymous):

no parentheses

OpenStudy (amistre64):

that just leaves order of operations i guess :)

OpenStudy (amistre64):

x - (4/x^3) - 9 +7x -18

OpenStudy (anonymous):

@amistre64 yes

OpenStudy (amistre64):

8x -27 -(4/x^3) 8x^4 -27x^3 -4 ---------------- x^3

OpenStudy (anonymous):

- (4/x^3) +8x -27

OpenStudy (anonymous):

x squared divided by x cubed - 9x square ___________ _____________________ x squared - 4 x squared + 7x-18

OpenStudy (anonymous):

sorry this is the question: x cubed divided by x cubed - 9x squared ___________ ______________________ x squared - 4 x squared +7x - 18

OpenStudy (amistre64):

\[\frac{\frac{x^3}{x^2-4}}{\frac{x^3 -9x^2}{x^2 +7x -18}} ???\]

OpenStudy (amistre64):

that came out tiny ;)

OpenStudy (anonymous):

thats it

OpenStudy (amistre64):

factor everything so you know what to cancel out; Also, take the bottom "fraction"; turn it over, spank it, and multiply it to the top "fraction"; like this: xxx (x+9)(x-2) --------- * ---------- (x-2)(x+2) xx (x-9)

OpenStudy (amistre64):

Now cancel like terms top tp bottom x(x+9) ---------- (x+2) (x-9)

OpenStudy (anonymous):

thank you, I had it right, I really appreciate it.

OpenStudy (amistre64):

youre welcome :) its alot easier to solve when you can "read" the problem ;)

OpenStudy (anonymous):

are you ready for another one

OpenStudy (amistre64):

I spose.... whatcha got?

OpenStudy (anonymous):

m squared - n squared -------------------- 5m - n m = - 1/5 and n= 1/5

OpenStudy (amistre64):

so you just fill in the variables with the values?

OpenStudy (amistre64):

1/25 - 1/25 = 0 0/anything = 0

OpenStudy (anonymous):

could someone please help me with factoring